The length of the base of an isosceles triangle is 39.01 inches. Each base angle is 23.44 degrees. find the length of each of the two equal sides of the triangle.
the two sides of the triangle have lengths of __ inches.

Answer by lwsshak3(11628)

The length of the base of an isosceles triangle is 39.01 inches. Each base angle is 23.44 degrees. find the length of each of the two equal sides of the triangle.
***
let x=length of each of rthe two equal sides
Draw a perpendicular line from the apex to the base, forming a right triangle where the adjacent side = base/2 and hypotenuse=one of the two equal sides,x)
cos 23.44˚=(39.01/2)/x
x=(39.01/2)/cos 23.44˚
x=21.26 inches